OpenStudy (anonymous):

Solve the following equations for all solutions of x 2sin^2x+3cos-3=0

hartnn (hartnn):

u need to convert sin^2 x into cos^2 x which i dentity u know of, does this ?

hartnn (hartnn):

an identity relating sin^2 x and cos^2 x is ?

OpenStudy (anonymous):

Pythagorean?

hartnn (hartnn):

yes, state that or find sin^2 x from that, what u get ?

OpenStudy (anonymous):

sin^2x+cos^2x=1

hartnn (hartnn):

yes so sin^2 x from ^ will be ?

OpenStudy (anonymous):

1+cos^2x

hartnn (hartnn):

sin^2x+cos^2x=1 so sin^2 x = 1-cos^2 x got this ??

OpenStudy (anonymous):

Oh...yeah, yeah, yeah

hartnn (hartnn):

now put that in 2sin^2x+3cos-3=0 and u get an equation only in terms of cos x

OpenStudy (anonymous):

2cos^2x+3cos-1=0?

hartnn (hartnn):

2sin^2x+3cos-3=0 2-2cos^2x +3cos x -3 = 0 -2cos^2 x + 3cos x -1 = 0 2cos^2 x -3cos x +1 = 0 got this ?

OpenStudy (anonymous):

Yes...now I have 2cos^x-3cosx+1=0

hartnn (hartnn):

now put cos x =y so u get a quadratic in y can u solve quadratic equations ?

OpenStudy (anonymous):

Just with a=2,b=-3,c=1?

hartnn (hartnn):

yes :) so u know the formula ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So is there going to be two answers then?

hartnn (hartnn):

what 2 answers u got ?

OpenStudy (anonymous):

I came up with x=3+-sqrt72/4

OpenStudy (anonymous):

Unsure though...

hartnn (hartnn):

72? b^2-4ac = (-3)^2 - 4(2)(1) = 9-8 = ?

OpenStudy (anonymous):

Ugh....I multiplied x4ac So...9-8=1

hartnn (hartnn):

yes, so two values of y are >?

OpenStudy (anonymous):

Is it 3+-1/4?

hartnn (hartnn):

so 2 values are ?

OpenStudy (anonymous):

13/4,11/4?

hartnn (hartnn):

13 ? 11 ? (3+1)/4 or (3-1) /4 so, 4/4 or 2/4 so, 1 or 1/2 ok?

OpenStudy (anonymous):

OMG...sorry. I was taking 3+1/4 It's all over 4, yes

hartnn (hartnn):

so u have cos x =1 or cos x = 1/2 right ? can u solve this 2 equations ?

OpenStudy (anonymous):

So cosign of 1 is pi/2

OpenStudy (anonymous):

And pi/6?

hartnn (hartnn):

cos (0) = 1 cos (pi)= 1 so cos x =1 will give x= 0,pi,2pi,....n*pi

OpenStudy (anonymous):

Thank you so much for your time. These processes are very hard for me. I know practice will help. Thank you again for walking me through each step.

hartnn (hartnn):

still cos x =1/2 is remaining

OpenStudy (anonymous):

So if cosx=1/2 then @ pi/3?

hartnn (hartnn):

yes, and other values ?

OpenStudy (anonymous):

Also 5pi/3 @ 300deg

OpenStudy (anonymous):

Thank goodness for a unit circle :)

hartnn (hartnn):

yeah, in general, it will be pi/3 + 2n*pi x= pi/3 , pi/3+ 2pi, pi/3 + 4pi, ...pi/3 + 2n*pi